# Less but Better: Enabling Generalized Zero-shot Learning Towards Unseen Domains by Intrinsic Learning from Redundant LLM Semantics

### v4

@@ Abstract
Generalized zero-shot learning (GZSL) focuses on recognizing seen and unseen classes against domain shift problem where data of unseen classes may be misclassified as seen classes. However, existing GZSL is still limited to seen domains. In the current work, we pioneer cross-domain GZSL (CDGZSL) which addresses GZSL towards unseen domains. Different from existing GZSL methods which alleviate domain shift problem by generating features of unseen classes with semantics, CDGZSL needs to construct a common feature space across domains and acquire the corresponding intrinsic semantics shared among domains to transfer from seen to unseen domains. Considering the information asymmetry problem caused by redundant class semantics annotated with large language models (LLMs), we present Meta Domain Alignment Semantic Refinement (MDASR). Technically, MDASR consists of two parts: Inter-class Similarity Alignment, which eliminates the non-intrinsic semantics not shared across all domains under the guidance of inter-class feature relationships, and Unseen-class Meta Generation, which preserves intrinsic semantics to maintain connectivity between seen and unseen classes by simulating feature generation. MDASR effectively aligns the redundant semantic space with the common feature space, mitigating the information asymmetry in CDGZSL. The effectiveness of MDASR is demonstrated on the Office-Home and Mini-DomainNet, and we have shared the LLM-based semantics for these datasets as the benchmark.
Index Terms:
@@
Generalized zero-shot learning, unseen domain, semantic refinement, large language model, information asymmetry.
I Introduction
@@
The era of Large Language Models (LLMs) marks a tremendous leap in the ability of artificial intelligence to understand and generate natural language. With massive numbers of parameters and extensive training data (for example, GPT-3 has 175 billion parameters and 753.4 GB of training data [1]), LLMs have the capacity to internalize vast amounts of knowledge and externalize it for application in the virtual and real world [2, 3]. In some machine learning studies that fuse data and knowledge [4, 5, 6], researchers previously relied on the knowledge from human experts to provide models with prior information, aiming to obtain results that outperform purely data-driven techniques. The emergence of LLMs has, to some extent, supplanted the role of human experts, thereby bringing promising prospects to data-knowledge fusion studies, such as zero-shot learning (ZSL) [7].
@@
ZSL is designed to recognize unseen classes not present in the training set by leveraging semantic knowledge to bridge the gap between seen and unseen classes, mimicking human cognitive inference [8, 9, 10]. Initially, ZSL approaches like DAP [9] and ALE [11] construct mappings from data to semantics, training on seen classes and extrapolating to unseen test data, under the assumption that only unseen classes would be tested. This works well when the assumption holds, but real-world scenarios often involve a mix of seen and unseen classes. To overcome this limitation, Generalized Zero-Shot Learning (GZSL) [12, 13, 14, 15] is introduced, requiring models to identify both seen and unseen classes during testing. However, ZSL models tend to misclassify unseen classes as seen due to the domain shift problem [16, 17], leading to poor GZSL performance. To combat this, generative methods such as f-CLSWGAN [13] and ABAGAN [18] have been developed, modeling between the semantic space and the feature space, generating unseen class data based on semantic knowledge to reduce model overfitting on seen classes and mitigate domain shift problem.
@@
Before the era of LLMs, ZSL and GZSL acquired semantic knowledge through the method of expert-annotated attributes [19]. This approach is only feasible for datasets with a relatively small number of classes due to the high cost of annotation and the reliance on expert experience. For large-scale datasets, researchers often resorted to using word vectors as a means of automatically acquiring knowledge of different classes [19]. However, the semantic information contained within these vectors is not rich enough, leading to potentially suboptimal experimental results. Fortunately, as mentioned earlier, the arrival of LLMs has provided a more efficient way to acquire knowledge [20]. By crafting appropriate prompts, we can extract a vast amount of descriptive text for each class from an LLM, and the rich semantic information contained in these texts allows for the performance of ZSL at a lower cost.
@@
Recently, the exploration of data from unseen domains and unseen classes simultaneously has garnered considerable attentions, which is called cross-domain zero-shot learning (CDZSL) [21]. The training set for CDZSL comprises data from seen classes across multiple seen domains, while the testing set consists of data from unseen classes within the unseen domain. Undoubtedly, because of the intervention of domains, CDZSL presents a greater challenge than ZSL due to its demand for a heightened level of model generalization. To address the gap of unseen classes and domains, Mancini et al. employed the Mixup technique [22] and introduced CuMix [21]. This approach utilizes the underlying structure of ZSL methods, mapping data to the feature space before subsequently mapping features to the semantic space. During the training process, samples from diverse domains and classes are intermixed within the data and feature spaces, effectively simulating novel domains and classes. Building upon CuMix, SEIC [23] introduced intermediate classes to imbue the feature embedding space with semantic meaning. The semantic knowledge utilized by these methods is still in the form of word vectors.
@@
Just like in ZSL, CDZSL also makes an assumption about the inference process: during the testing phase, only unseen classes from unseen domains are present. We believe that the assumption of CDZSL does not agree well with reality, as in real-world scenarios, both seen and unseen classes are likely to occur. Hence, we first introduce Cross-Domain Generalized Zero-Shot Learning (CDGZSL) and demonstrate it in Fig. 1. In CDGZSL, the training set includes seen class data from multiple seen domains, while the test set includes both seen and unseen class data from the unseen domain. A comparison between zero-shot learning, generalized zero-shot learning, cross-domain zero-shot learning, and cross-domain generalized zero-shot learning can be seen in Table I. Existing studies encompass zero-shot learning, generalized zero-shot learning and cross-domain zero-shot learning. Generalized zero-shot learning removes the inference assumption of zero-shot learning but maintains the assumption of the single domain. In contrast, cross-domain zero-shot learning, despite its ability to handle multiple domains, does not eliminate the inference assumption. This demonstrates that generalized zero-shot learning and cross-domain zero-shot learning both have assumptions, and cross-domain generalized zero-shot learning breaks the limitations of both, resulting in test scenarios that agree well with reality.
@@
TABLE I: Comparison among several ZSL settings.
@@
    indicates the presence of research or assumption, while ✘represents the opposite meaning.
Although generative methods could be employed to address the transition from zero-shot learning to generalized zero-shot learning, the same approach cannot be directly applied from cross-domain zero-shot learning to cross-domain generalized zero-shot learning. For the sake of cost-effectiveness and speed, a large volume of class description text can be obtained from LLMs as semantics, and is applied to all domains (including seen and unseen domains). However, this leads to the asymmetry between the low information density of semantic modality and the high information density of feature modality. In terms of the semantic modality, a significant portion of the large-volume LLM semantics is not intrinsic semantics which are shared across domains, leading to a low knowledge information density. For the feature modality, in order to generalize from seen domains to unseen domains, the generator needs to produce features within a common feature space shared by all domains. However, this common feature space encapsulates the essence of each class and thus has a higher information density. This asymmetry poses a challenge for the generator to accurately model the relationship between these two modalities. It results in greater differences between the generated features for unseen classes and their actual features, making it difficult to address the domain shift problem when transitioning from CDZSL to CDGZSL.
In this paper, we introduce a novel approach named Meta Domain Alignment Semantic Refinement (MDASR) aimed at enabling GZSL towards unseen domains. To address the aforementioned issue of information asymmetry, unlike existing ZSL, GZSL and CDZSL endeavors, we learn intrinsic information from the redundant semantics and perform generative modeling between the refined semantics and domain common features, which are symmetrical in information density, to mitigate the domain shift problem. To this end, MDASR includes two optimization objectives: Inter-class Similarity Alignment (ISA) and Unseen-class Meta Generation (UMG). Guided by the feature space, ISA eliminates non-intrinsic semantics while UMG preserves intrinsic semantics. Specifically, on the one hand, ISA leverages the fundamental and invariant relationships between classes, which are derived from multi-domain data, as a standard to remove non-intrinsic semantics. On the other hand, UMG simulates the generation process of unseen class features to ensure the semantic linkage from seen to unseen classes remains unbroken, thus preserving intrinsic semantics. In summary, our method, through ISA and UMG, addresses the issue of information asymmetry between semantic and feature spaces, thereby alleviating the domain shift problem in CDGZSL. The total contributions of this paper are summarized below.
@@
    We first verify the CDGZSL capability of MDASR on two open-source datasets, Office-Home and Mini-DomainNet, and we have made the corresponding LLM-based semantic texts for these datasets available to the public, providing a benchmark for subsequent researchers.
II Motivation
In this section, we would like to clarify our motivation: the problem of semantic variation and redundancy, and the resulting information asymmetry between semantic space and feature space that makes the solving of domain shift problem in CDGZSL more difficult.
II-A The problem of semantic variation and redundancy
We would like to introduce the semantic variation caused by multiple domains, as well as the issue of redundancy arising from the general semantic descriptions.
II-A1 Semantic variation
@@ II-A2 Semantic redundancy
II-B Information asymmetry between semantics and features
@@
If we would like to address domain shift problem, a generator is needed to model the relationship between the modalities of semantics and features and generate features of unseen classes [24]. However, there is an information asymmetry between the two modalities, as shown in Fig. 5, impeding the effectiveness of generative methods in transitioning from CDZSL to CDGZSL.
The information asymmetry between the semantic and feature modalities originates from their different information densities. In this paper, information density is defined as the proportion of shared information across domains relative to all information. For the semantic modality, the intrinsic semantics shared by each domain only account for a portion of the general semantics provided by the LLM, while the rest of the information is non-intrinsic, resulting in low information density, as shown in the left half of Fig. 5. In contrast, for the feature modality, existing transfer learning methods can extract the common feature space across domains, eliminating domain-specific information, thus leading to high information density, as represented by the dark blue overlapping area in the right half of Fig. 5.
In fact, the solution lies in obtaining the common features and identifying the corresponding intrinsic information within the semantics. The prior problem has been resolved by predecessors, making it simpler, whereas the latter remains unsolved, presenting more of a challenge. If the information asymmetry is not tackled, the generator could only be trained with non-intrinsic semantics mixed together, which harms the address of domain shift problem.
@@
III Methodology
In this section, we first introduce the problem formulation and notations used in this paper. From there, the process of utilizing LLMs for extracting semantics of various classes is explored. This is followed by a recapitulative explanation of the proposed MDASR. Then, the detailed paradigm for visual alignment and meta semantic refinement is presented. We end the section with the CDGZSL inference procedure.
III-A Problem formulation and notations
In the task of CDGZSL, our goal is to recognize both seen and unseen classes in the unseen domain. Only labeled seen class samples from multiple seen domains and their corresponding semantics are provided during training.
@@ III-A Problem formulation and notations
Given $\mathcal{S}$, we aim at learning a function mapping an image $x$ of unseen domains $\mathcal{D}^{u}\in\mathcal{D}$ to its corresponding label in a set of classes $\mathcal{Y}$. $\mathcal{D}^{s}\cap\mathcal{D}^{u}=\varnothing$. Following the setting of GZSL, $\mathcal{Y}=\mathcal{Y}^{s}\cup\mathcal{Y}^{u}$, where $\mathcal{Y}^{u}$ are the labels space of unseen classes, and $\mathcal{Y}^{s}\cap\mathcal{Y}^{u}=\varnothing$. Note that during testing, $\mathcal{A}^{u}$ are provided, which is the semantic space of unseen classes.
III-B Extracting semantics via LLMs with prompt engineering
@@
III-C Overall framework of MDASR
@@
To achieve CDGZSL in $\Psi$, we need to generate reliable features of unseen classes based on their semantics. As discussed in Section II, it is crucial to obtain the intrinsic semantics to mitigate the issues of semantic redundancy. Consequently, we introduce a meta semantic refinement framework designed to optimize the semantic encoder $E_{s}$, guided by features within $\Psi$. This approach preserves valuable semantic information for future feature generation. Finally, by utilizing the generated unseen features in conjunction with real seen features, we can facilitate cross-domain generalized zero-shot classification of samples from unseen domains.
III-D Adversarial construction of common feature space
To construct the common feature space across different domains, we employ the Domain-Adversarial Neural Network technique [25]. For this part, specifically refer to the upper half of Fig. 6. We establish three networks: the data encoder $E_{d}$, the domain classifier $C_{d}$, and the seen-class classifier $C_{s}$, parameterized by $\theta$, $\mu$, and $\sigma$ respectively. It should be noted that only training data from seen classes in the seen domains is provided.

### v5

# Enabling Generalized Zero-shot Learning Towards Unseen Domains by Intrinsic Learning from Redundant LLM Semantics

@@ Abstract
Generalized zero-shot learning (GZSL) focuses on recognizing seen and unseen classes against domain shift problem where data of unseen classes may be misclassified as seen classes. However, existing GZSL is still limited to seen domains. In the current work, we study cross-domain GZSL (CDGZSL) which addresses GZSL towards unseen domains. Different from existing GZSL methods, CDGZSL constructs a common feature space across domains and acquires the corresponding intrinsic semantics shared among domains to transfer from seen to unseen domains. Considering the information asymmetry problem caused by redundant class semantics annotated with large language models (LLMs), we present Meta Domain Alignment Semantic Refinement (MDASR). Technically, MDASR consists of two parts: Inter-class similarity alignment, which eliminates the non-intrinsic semantics not shared across all domains under the guidance of inter-class feature relationships, and unseen-class meta generation, which preserves intrinsic semantics to maintain connectivity between seen and unseen classes by simulating feature generation. MDASR effectively aligns the redundant semantic space with the common feature space, mitigating the information asymmetry in CDGZSL. The effectiveness of MDASR is demonstrated on two datasets, Office-Home and Mini-DomainNet, and we have shared the LLM-based semantics for these datasets as a benchmark.
keywords:
@@
Generalized zero-shot learning , unseen domain , semantic refinement , large language model , information asymmetry
1 Introduction
@@
The era of Large Language Models (LLMs) marks a tremendous leap in the ability of artificial intelligence to understand and generate natural language. With massive parameters and extensive training data, LLMs have the capacity to internalize vast amounts of knowledge and externalize it for application in the virtual and real world (Zhao and Zhou, 2023; Chang et al., 2024). In some machine learning studies that fuse data and knowledge (Pan et al., 2025; Li and Zhao, 2023), researchers previously relied on the knowledge from human experts to provide models with prior information, aiming to obtain results that outperform purely data-driven techniques (Lu et al., 2025; Zhao, 2022). The emerging LLMs have, to some extent, supplanted the role of human experts, thereby bringing promising prospects to data-knowledge fusion studies, such as zero-shot learning (ZSL) (Wang et al., 2020; Gautam et al., 2022; Song et al., 2024).
@@
ZSL is designed to recognize unseen classes not present in the training set by leveraging semantic knowledge to bridge the gap between seen and unseen classes, mimicking human cognitive inference (Feng and Zhao, 2021; Lampert et al., 2013; Chen et al., 2024). Initially, ZSL approaches like DAP (Lampert et al., 2013) and ALE (Akata et al., 2015) construct mappings from data to semantics, training on seen classes and extrapolating to unseen test data, under the assumption that only unseen classes would be tested. This works well when the assumption holds, but real-world scenarios often involve a mix of seen and unseen classes. To overcome this limitation, Generalized Zero-Shot Learning (GZSL) (Feng et al., 2022; Zhang and Li, 2025) is introduced, requiring models to identify both seen and unseen classes during testing. However, ZSL models tend to misclassify unseen classes as seen due to the domain shift problem (Ye et al., 2022; Yue et al., 2024), leading to poor GZSL performance. To combat this, generative methods such as ZeroDiff (Ye et al., 2025) and VADS (Hou et al., 2024) have been developed, modeling between the semantic space and the feature space, generating unseen class data based on semantic knowledge to reduce model overfitting on seen classes and mitigate domain shift problem.
@@
Before the era of LLMs, ZSL and GZSL acquired semantic knowledge through the method of expert-annotated attributes (Sun et al., 2021; Rossi et al., 2024). This approach is only feasible for datasets with a relatively small number of classes due to the high cost of annotation. Some professional fields even require seeking knowledge from experts (Feng and Zhao, 2020). For large-scale datasets, researchers often resorted to using word vectors as a means of automatically acquiring knowledge of different classes (Sun et al., 2021). However, the semantic information contained within these vectors is not rich enough, leading to potentially suboptimal experimental results. Fortunately, as mentioned earlier, the arrival of LLMs has provided a more efficient way to acquire knowledge (Naeem et al., 2023). By crafting appropriate prompts, we can extract a vast amount of descriptive text for each class from an LLM, and the rich semantic information contained in these texts allows for improving the performance of ZSL at a lower cost.
@@
Recently, the simultaneous exploration of data from unseen domains and unseen classes has garnered considerable attentions, which is called cross-domain zero-shot learning (CDZSL) (Mancini et al., 2020). The training set for CDZSL comprises data from seen classes across multiple seen domains, while the testing set consists of data from unseen classes within the unseen domain. Undoubtedly, because of the intervention of domains, CDZSL presents a greater challenge than ZSL due to its demand for a heightened level of model generalization. To address the gap of unseen classes and domains, Mancini et al. employed the Mixup technique (Zhang et al., 2018) and introduced CuMix (Mancini et al., 2020). This approach utilizes the underlying structure of ZSL methods, mapping data to the feature space before subsequently mapping features to the semantic space. During the training process, samples from diverse domains and classes are intermixed within the data and feature spaces, effectively simulating novel domains and classes. Building upon CuMix, SEIC (Mondal and Biswas, 2022) introduced intermediate classes to add the feature embedding space with semantic meaning.
@@
Just like in ZSL, CDZSL also makes an assumption about the inference process: during the testing phase, only unseen classes from unseen domains are present. We believe that the assumption of CDZSL does not agree well with reality, as in real-world scenarios, both seen and unseen classes are likely to occur. Hence, we first introduce Cross-Domain Generalized Zero-Shot Learning (CDGZSL) and demonstrate it in Fig. 1. In CDGZSL, the training set includes seen class data from multiple seen domains, while the test set includes both seen and unseen class data from the unseen domain. A comparison between zero-shot learning, generalized zero-shot learning, cross-domain zero-shot learning, and cross-domain generalized zero-shot learning can be seen in Table 1. Existing studies encompass zero-shot learning, generalized zero-shot learning and cross-domain zero-shot learning. Generalized zero-shot learning removes the inference assumption of zero-shot learning but maintains the assumption of the single domain. In contrast, cross-domain zero-shot learning, despite its ability to handle multiple domains, does not eliminate the inference assumption. This demonstrates that generalized zero-shot learning and cross-domain zero-shot learning both have assumptions, and cross-domain generalized zero-shot learning breaks the limitations of both, resulting in test scenarios that agree well with reality.
@@
Table 1: Comparison among several ZSL settings.
@@
    ✓indicates the presence of research or assumption, while ✗represents the opposite meaning.
Although generative methods could be employed to address the transition from zero-shot learning to generalized zero-shot learning, the same approach cannot be directly applied from cross-domain zero-shot learning to cross-domain generalized zero-shot learning. As mentioned earlier, a large volume of class description text can be obtained from LLMs as semantics for the sake of cost-effectiveness and speed, and is applied to all domains (including seen and unseen domains). However, this leads to the asymmetry between the low information density of semantic modality and the high information density of feature modality. In terms of the semantic modality, a significant portion of the large-volume LLM semantics is not intrinsic semantics which are shared across domains, leading to a low knowledge information density. For the feature modality, in order to generalize from seen domains to unseen domains, the generator needs to produce features within a common feature space shared by all domains. However, this common feature space encapsulates the essence of each class and thus has a higher information density. This asymmetry poses a challenge for the generator to accurately model the relationship between these two modalities. It results in greater differences between the generated features for unseen classes and their actual features, making it difficult to address the domain shift problem when transitioning from CDZSL to CDGZSL.
In this paper, we introduce a novel approach named Meta Domain Alignment Semantic Refinement (MDASR) aimed at enabling GZSL towards unseen domains. To address the aforementioned issue of information asymmetry, unlike what the existing ZSL, GZSL and CDZSL endeavors to, we learn intrinsic information from the redundant semantics and perform generative modeling between the refined semantics and domain common features, which are symmetrical in information density, to mitigate the domain shift problem. To this end, MDASR includes two optimization objectives: Inter-class Similarity Alignment (ISA) and Unseen-class Meta Generation (UMG). Guided by the feature space, ISA eliminates non-intrinsic semantics while UMG preserves intrinsic semantics. Specifically, on the one hand, ISA leverages the fundamental and invariant relationships between classes, which are derived from multi-domain data, as a standard to remove non-intrinsic semantics. On the other hand, UMG simulates the generation process of unseen class features to ensure that the semantic linkage from seen to unseen classes remains unbroken, thus preserving intrinsic semantics. In summary, our method, through ISA and UMG, addresses the issue of information asymmetry between semantic and feature spaces, thereby alleviating the domain shift problem in CDGZSL. The total contributions of this paper are summarized below.
@@
    We have verified the CDGZSL capability of MDASR on two open-source datasets, Office-Home and Mini-DomainNet, and we have then made the corresponding LLM-based semantic texts for these datasets available to the public, providing a benchmark for interested researchers.
2 Motivation
In this section, we will clarify our motivation: the problem of semantic variation and redundancy, and the resulting information asymmetry between semantic space and feature space that makes the solution of domain shift problem in CDGZSL challenging.
2.1 The problem of semantic variation and redundancy
We introduce the semantic variation caused by multiple domains, as well as the issue of redundancy arising from the general semantic descriptions.
2.1.1 Semantic variation
@@ 2.1.2 Semantic redundancy
2.2 Information asymmetry between semantics and features
@@
If we want to address domain shift problem, a generator is needed to model the relationship between the modalities of semantics and features, and generate features of unseen classes (Pourpanah et al., 2022). However, there is an information asymmetry between the two modalities, as shown in Fig. 5, impeding the effectiveness of generative methods in transitioning from CDZSL to CDGZSL.
The information asymmetry between the semantic and feature modalities originates from their different information densities. In this paper, information density is defined as the proportion of shared information across domains relative to all information. For the semantic modality, the intrinsic semantics shared by each domain only account for a portion of the general semantics provided by the LLM, while the rest of the information is non-intrinsic, resulting in low information density, as shown in the left half of Fig. 5. In contrast, for the feature modality, domain adaptation methods can extract the common feature space across domains, eliminating domain-specific information, thus leading to high information density, as represented by the dark blue overlapping area in the right half of Fig. 5.
Existing approaches to enhance model transferability include transfer learning (Ganin et al., 2016; Lu et al., 2025) and meta-learning (Finn et al., 2017; Nichol et al., 2018). However, these methods face significant limitations in addressing the information asymmetry caused by semantic redundancy. Transfer learning focuses on feature alignment but ignores the redundancy in semantic annotations. Meta-learning optimizes for task-specific adaptation but does not consider cross-domain semantic variations. The previously mentioned zero-shot learning, while leveraging semantic knowledge, assumes that semantics are domain-agnostic, failing to handle domain-specific redundancy. As summarized in Table 2, none of these methods explicitly address the information asymmetry between semantic and feature modalities, which is critical for cross-domain generalized zero-shot learning (CDGZSL).
Paradigm Representative Works Problem Solved Consider Information Asymmetry? Transfer Learning $\cdot$ DANN (Ganin et al., 2016) $\cdot$ MAN (Lu et al., 2025) Align different domains Meta-Learning $\cdot$ MAML (Finn et al., 2017) $\cdot$ Reptile (Nichol et al., 2018) Fast adaptation to new tasks Existing ZSL $\cdot$ DAP (Lampert et al., 2013) $\cdot$ CuMix (Mancini et al., 2020) $\cdot$ Zerodiff (Ye et al., 2025) Recognize unseen classes CDGZSL $\cdot$ MDASR (Proposed) Refine redundant semantics, thus recognize seen & unseen classes in useen domain
Table 2: Comparison between existing paradigms and the proposed MDASR method. ✓indicates that the method consider information asymmetry, while ✗represents the opposite.
As shown in Table 2, existing paradigms cannot be simply combined to solve the problem of information asymmetry in CDGZSL. Therefore, a novel approach is required to explicitly address this problem. Our proposed MDASR achieves this by refining redundant semantics and preserving intrinsic ones, thereby mitigating information asymmetry and enabling effective CDGZSL. We will explain our method in the following section.
@@
3 Methodology
In this section, we first introduce the problem formulation and notations used in this paper. From there, the process of utilizing LLMs for extracting semantics of various classes is explored. This is followed by a recapitulative explanation of the proposed MDASR. Then, the detailed paradigm for visual alignment and meta semantic refinement is presented. We end the section with introducing the CDGZSL inference procedure.
3.1 Problem formulation and notations
In the task of CDGZSL, our goal is to recognize both seen and unseen classes in the unseen domain. Only labeled seen class samples from multiple seen domains and their corresponding semantics are provided during the training.
@@ 3.1 Problem formulation and notations
Given $\mathcal{S}$, we aim at learning a function mapping an image $x$ of unseen domains $\mathcal{D}^{u}\in\mathcal{D}$ to its corresponding label in a set of classes $\mathcal{Y}$. $\mathcal{D}^{s}\cap\mathcal{D}^{u}=\varnothing$. Following the setting of GZSL, $\mathcal{Y}=\mathcal{Y}^{s}\cup\mathcal{Y}^{u}$, where $\mathcal{Y}^{u}$ are the labels space of unseen classes, and $\mathcal{Y}^{s}\cap\mathcal{Y}^{u}=\varnothing$. Note that during the testing, $\mathcal{A}^{u}$ are provided, which is the semantic space of unseen classes.
3.2 Extracting semantics via LLMs with prompt engineering
@@
3.3 Overall framework of MDASR
@@
To achieve CDGZSL in $\Psi$, we need to generate reliable features of unseen classes based on their semantics. As discussed in Section 2, it is crucial to obtain the intrinsic semantics to mitigate the issues of semantic redundancy. Consequently, we introduce a meta semantic refinement framework designed to optimize the semantic encoder $E_{s}$, guided by features within $\Psi$. This approach preserves valuable semantic information for future feature generation. Finally, by utilizing the generated unseen features in conjunction with real seen features, we can facilitate cross-domain generalized zero-shot classification of samples from unseen domains.
3.4 Adversarial construction of common feature space
To construct the common feature space across different domains, we employ the Domain-Adversarial Neural Network technique (Ganin et al., 2016). For this part, specifically refer to the upper half of Fig. 6. We establish three networks: the data encoder $E_{d}$, the domain classifier $C_{d}$, and the seen-class classifier $C_{s}$, parameterized by $\theta$, $\mu$, and $\sigma$ respectively. It should be noted that only training data from seen classes in the seen domains is provided.